# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

### v1

@@ 3.4 Anisotropy of emergent large-scale dynamics
While the previous section 3.3 has made clear that vertical aspects of the flow dynamics converge for horizontally highly extended domains only, section 3.1 has demonstrated that the forced shear flows considered here develop certain characteristic horizontal structures, such as the primary Kelvin-Helmholtz billows. Therefore, it is appropriate to investigate whether there are emergent horizontally-aligned structures in the statistically steady flow at later times (when such early billows have broken down).
@@
Figure 6 visualises both the early and final emergent dynamics present in our largest square domain of $L_{\mathrm{h}}=512$. Panels (a – d, i – l) depict the entire horizontal cross-section at midplane, $\Phi\left(z=0\right)$, whereas panels (e – h, m – p) depict an associated vertical slice at $\Phi\left(y=0\right)$. Note that the vertical slices of $b$ from panels (h, p) remind of figure 2 (c, e) despite the domain now being $16$ times as large. A video of the evolution of the flow in this domain of $L_{\mathrm{h}}=512$ – from $t=0$ to $t=t_{\textrm{evo}}$ – is provided as supplemental material.
At the earlier time, as shown in panel (h), the flow is clearly associated with the growth of the initial KHI and the subsequent formation of overturning billows, analogously to the simulation discussed in section 3.1 and shown in figure 2. Interestingly, as becomes clear via the associated horizontal slice in panel (d), although the initial overturning billows may extend across the entire extended spanwise direction, knots, tubes and defects between these rolls may introduce defects and imperfections to this otherwise regular pattern, as previously observed experimentally (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and numerically (Fritts et al., 2022a, b) for (in the spanwise direction) sufficiently wide domains. However, these early aspects of the flow dynamics appear (at least superficially) to be absent or smeared out by the sustained turbulence at late times as shown in panels (l, p).
@@
Independently of the specific point in time, this comparison of vertical slices with the corresponding horizontal midplanes demonstrates the co-existence of apparently quasi-regular dynamics in both the streamwise and spanwise directions. However, these visualisations suggest that the horizontal structures have characteristic length scales much larger than the shear layer (half-)depth, i.e., $\left\{d,\delta\right\}\ll\Lambda_{\textrm{hor}}$, implying a strong scale separation or anisotropy in the flow.
@@
of various quantities. Here, $\hat{\Phi}$ represent the Fourier coefficients corresponding to a transform in either the streamwise or spanwise direction, the asterisk denotes the complex conjugate, and $k_{x}$ and $k_{y}$ are the associated streamwise and spanwise components of the wave vector, respectively. In order to allow for a direct comparability of these spectra with their corresponding term in the kinetic energy or buoyancy variance equation – see equations (26) and (27), respectively, in appendix A – via Parseval’s theorem, the coefficient $C$ depends on the particular quantities $\Phi_{1}$ and $\Phi_{2}$: $C=1/2$ for $\Phi_{1}=\Phi_{2}\in\left\{u_{x},u_{y},u_{z},b\right\}$, $C=\mathrm{Ri}_{0}$ for $\Phi_{1}=u_{z}\wedge\Phi_{2}=b$ (or vice versa), $C=2/\mathrm{Re}_{0}$ for $\Phi_{1}=\Phi_{2}=\bm{S}$, or $C=1/\left(\mathrm{Re}_{0}\mathrm{Pr}\right)$ for $\Phi_{1}=\Phi_{2}=\nabla b$. We summarise the key results in figure 7.
@@
Figure 7: Magnitude quantification of emergent structures. (a) Emergent flow structures offer pronounced spectral peaks $\hat{\lambda}$ in the streamwise directions for all variables $\left\{u_{x},u_{y},u_{z},b\right\}$. In contrast, (b) only the streamwise velocity $u_{x}$ exhibits a pronounced peak in spanwise direction. (c) A systematic comparison of characteristic extensions of emergent flow structures along the streamwise, spanwise, and vertical directions highlights a strong anisotropy in large-scale dynamics. Note that while the coloured lines in panels (a, b) correspond to our largest square domain with $L_{\mathrm{h}}=512$, grey lines are extracted from even more extended but non-square domains ($L_{x}=2048$ and $L_{y}=512$ or $L_{y}=2048$ and $L_{x}=512$). Moreover, as the dash-dotted line $L_{x}=L_{y}$ in panel (c) demonstrates that flow structures may clearly be limited or affected by horizontally insufficiently extended domains, only horizontal extents $L_{\mathrm{h}}\gtrsim 256$ are large enough to resolve the most extended flow structures in the horizontal direction.
@@
In panel (a), we plot the streamwise energy spectra of $\left\{u_{x},u_{y},u_{z},b\right\}$, subject to appropriate spatio-temporal averages. Remarkably, we find that all these flow fields exhibit pronounced spectral peaks at streamwise wave numbers $\hat{k}_{x}\sim\mathcal{O}\left(10^{-2}\right)$. This implies that a certain (narrow range of) streamwise wavelength(s) $\hat{\lambda}_{x}=2\pi/\hat{k}_{x}$ is particularly energetic. In other words, these pronounced spectral peaks establish the existence of flow structures with a characteristic streamwise length scale $\Lambda_{x}\sim\mathcal{O}\left(10^{2}\right)$ in all of these flow fields, thus implying that there is a preferred length scale for the emergent streamwise self-organisation of the large-scale dynamics. We propose a potential dynamical origin of this characteristic scale $\Lambda_{x}$ in section 4.
@@
As shown in panel (b), this emergent property of the streamwise spectra is in clear contrast to the behaviour of the spanwise spectra. On the one hand, we find a similarly pronounced spectral peak only for the $x$-component of the velocity field in the spanwise direction. Together with $\hat{k}_{y}\sim\mathcal{O}\left(10^{-1}\right)$, this implies again the existence of a characteristic spanwise length scale $\Lambda_{y}\sim\mathcal{O}\left(10^{1}\right)$. On the other hand, the other flow fields do not exhibit such a peak but rather flatten out at small $k_{y}$, demonstrating that there is no preferred characteristic length scale for the spanwise self-organisation of the large-scale dynamics in $\left\{u_{y},u_{z},b\right\}$. Note that even though $E_{u_{y}u_{y}}(k_{y,\textrm{min}})<E_{u_{y}u_{y}}(\hat{k}_{y})$, this difference is smaller than a factor of two and so we do not consider the associated maximum to be a ‘pronounced’ spectral peak.
@@
It is natural to ask whether there could be additional spectral peaks at even smaller wave numbers, i.e., whether our largest square domains of $L_{\mathrm{h}}=512$ might still be too small to capture additional emergent dynamics at even larger length scales. For this reason, we have conducted two additional simulations which increase either the streamwise or the spanwise extent of the domain by another factor of four. This results in $L_{x}\times L_{y}=2048\times 512$ or $512\times 2048$, respectively. We include the associated energy spectra in figure 7 (a, b) as grey solid lines. We find no evidence of such additional spectral peaks. Furthermore, as both the location and amplitude of the peaks from these spectra coincide with the ones from our largest square domain, these spectral peaks are shown to be characteristic of this particular flow and (crucially) independent of the horizontal extent of the domain. This implies that the large-scale dynamics is (at such large $L_{x}$ and $L_{y}$) indeed governed by mechanisms intrinsic to the flow. This is supported by appendix C, where we show that the energy spectra derived from smaller (yet sufficiently large) domains are also shown to converge with the present ones from the largest domains.
@@
$L_{x}\times L_{y}\thickspace\thickspace$ $E_{u_{x}u_{x}}$ $E_{u_{y}u_{y}}$ $E_{u_{z}u_{z}}$ $E_{bb}$ $E_{u_{z}b}$ $E_{u_{x}u_{x}}$ $2\times d$ $2\times\delta$
$16^{2}$ 00 $016\pm\hphantom{0}$ 0 $16\pm\hphantom{0}0$ $16\pm\hphantom{0}0$ $016\pm\hphantom{0}$ 0 $16\pm\hphantom{0}0$ $15\pm\hphantom{0}3$ 05.3±0.3plus-or-minus5.30.35.3$\pm$ 0.35.3 ± 0.3 05.6±0.3plus-or-minus5.60.35.6$\pm$ 0.35.6 ± 0.3
$32^{2}$ 00 $032\pm\hphantom{0}$ 0 $32\pm\hphantom{0}0$ $32\pm\hphantom{0}0$ $032\pm\hphantom{0}$ 0 $32\pm\hphantom{0}0$ $19\pm\hphantom{0}6$ $10.7\pm 0.7$ $11.1\pm 0.7$
$64^{2}$ 00 $064\pm\hphantom{0}$ 0 $64\pm\hphantom{0}0$ $64\pm\hphantom{0}2$ $064\pm\hphantom{0}$ 0 $64\pm\hphantom{0}0$ $33\pm 14$ $15.1\pm 0.9$ $15.5\pm 0.7$
@@
$128^{2}$ 00 $128\pm\hphantom{0}0$ $89\pm 32$ $77\pm 27$ $108\pm 30$ $89\pm 32$ $52\pm 24$ $16.1\pm 0.8$ $16.8\pm 0.9$
$256^{2}$ 00 $117\pm 18$ $84\pm 13$ $74\pm 12$ $92\pm 20$ $80\pm 12$ $52\pm 27$ $15.8\pm 0.4$ $16.4\pm 0.3$
$512^{2}$ 00 $108\pm 19$ $82\pm 10$ $76\pm\hphantom{0}9$ 089±11plus-or-minus891189$\pm$ 1189 ± 11 $79\pm\hphantom{0}8$ $49\pm 15$ $15.5\pm 0.2$ $16.1\pm 0.1$
$2048\times 512$ $111\pm 17$ $86\pm 12$ $77\pm 10$ 096±12plus-or-minus961296$\pm$ 1296 ± 12 $91\pm 10$ $46\pm 10$ $15.6\pm 0.1$ $16.2\pm 0.1$
$512\times 2048$ $118\pm 13$ $83\pm 10$ $75\pm\hphantom{0}8$ 095±11plus-or-minus951195$\pm$ 1195 ± 11 $86\pm 12$ $45\pm 15$ $15.5\pm 0.1$ $16.2\pm 0.1$
@@
Table 2: Emergent dynamical scales. This table quantifies the streamwise and spanwise or vertical extent of emergent large-scale dynamics via the wavelength associated with the spectral peak, the shear layer half-depth or the density interface half-depth. Values of the temporal mean and standard deviation are listed, the latter of which might be significant due to the discrete nature of wave numbers. Unreliable values, i.e. when the domains are insufficiently extended, are displayed in grey.
@@
Quantifying characteristic length scales associated with the emergent large-scale dynamics, we extract the wavelengths corresponding to these pronounced spectral peaks (or these spectra’s global maxima) for each simulation and summarise them in table 2. Additionally, figure 7 (c) visualises this dependence on $L_{\mathrm{h}}$. For small $L_{\mathrm{h}}$, as highlighted by the dash-dotted line, these extracted horizontal length scales $\Lambda_{\textrm{x, y}}$ are clearly biased by the domain size. However, this changes once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$ with the critical horizontal aspect ratio $L_{\mathrm{h,crit}}$ depending on the solution field (and, thus, the associated final characteristic length scale). While most of the peaks are already properly represented given $L_{\mathrm{h,crit}}=96$, some of them – such as $\hat{\lambda}_{x}\left(E_{u_{x}u_{x}}\right)$ – may require $L_{\mathrm{h}}=256$ to be fully resolved. Hence, a full convergence of the characteristic length scales associated with the large-scale dynamics is reached for $L_{\mathrm{h}}\gtrsim 256$ only.
@@
In addition to the characteristic streamwise and spanwise length scales associated with the large-scale dynamics, we include their characteristic vertical extent via the (total) shear layer depth in figure 7 (c). Note that since $d\approx\delta$ or $R\approx 1$, these data points equivalently show the (total) buoyancy interface depth. Comparing the characteristic length scales in the three spatial directions (as denoted by different symbol types), there is a clearly apparent anisotropy of the large-scale dynamics. While vertical scales $\Lambda_{\textrm{z}}\approx 16$ (circles) and spanwise scales $\Lambda_{\textrm{y}}\approx 50$ (upright triangles), streamwise scales (sideways triangles) may extend up to $\Lambda_{\textrm{x}}\approx 115$. Consequently, we find a hierarchy $\Lambda_{\textrm{z}}<\Lambda_{\textrm{y}}<\Lambda_{\textrm{x}}$ that spans across one order of magnitude.
@@
Having extracted these characteristic scales of large-scale dynamics, there are two particularly striking observations. Firstly, while $u_{z}$ and $b$ are related to the vertical transport of buoyancy via equation (16), their individual scales $\hat{\lambda}_{x}$ in $E_{u_{z}u_{z}}$ and $E_{bb}$ differ significantly. Secondly, the depth of the mixing zone, representing the characteristic vertical extent $\Lambda_{\textrm{z}}$, actually converges at smaller domain extents $L_{\mathrm{h}}$ compared with $\Lambda_{\textrm{x}}$ from $E_{bb}$. These two observations suggest the natural question as to why or how can the vertical characteristics (in both the velocity and buoyancy field) converge before the domain is sufficiently large to allow the separate flow fields to converge horizontally?
@@
Figure 8: Buoyancy exchange and dissipation. (a) The buoyancy flux (i.e. the vertical buoyancy advection) across the midplane – visualised via $B\left(x,y,z=0,t=t_{\textrm{evo}}\right)$ for $L_{\mathrm{h}}=512$ – generally stabilises the configuration. (b) The associated streamwise co-spectrum establishes the existence of a characteristic corresponding scale via a pronounced peak, which is in contrast to the spanwise direction. (c) Most of the dissipation, however, is associated with smaller scales similar to the mixing zone depth. In panels (b, c), coloured and grey spectra correspond – similar to figure 7 (a, b) – to the largest square and non-square domains, respectively.
@@
4 Discussion and perspective
Aiming to improve our fundamental understanding of the dynamics of forced stratified shear flows, motivated by their relevance to many geophysical flows, we have conducted a series of direct numerical simulations in computational domains of different sizes. Although our (forced) flow configuration is prone to primary Kelvin-Helmholtz instability, our configuration additionally and importantly includes a continuous forcing which tends to restore the initial profiles and ensures that the ensuing turbulence can be sustained for arbitrarily long times. Since it is this (at least quasi-) continuous forcing which connects to the motivating geophysical flows, we largely disregard the initial transient but rather focus on the statistically stationary dynamics during later time.
By varying the horizontal extent of the numerical domain from $16$ to $512$ times the initial shear layer (half-) depth $d_{0}$, we have demonstrated the emergence of strongly anisotropic large-scale dynamics. Provided a sufficiently large computational domain, the emergent characteristic length scales are $\Lambda_{z}\approx 16$ in the vertical direction, $\Lambda_{y}\approx 50$ in the spanwise direction, and up to $\Lambda_{x}\approx 115$ in the streamwise direction. In other words, we find an anisotropic hierarchy $\Lambda_{z}<\Lambda_{y}<\Lambda_{x}$ that spans across one order of magnitude. Crucially, the associated flow structures as well as their induced stirring, dissipation and irreversible mixing are remarkably sensitive to the confinement and start to become independent of the computational domain for extended domains of $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$ only. This convergence can be associated with the eventual resolution of the vertical buoyancy flux, which exhibits a peak characteristic streamwise scale $\Lambda_{x}\approx 85$. This scale represents approximately five times the final (total) depth of the turbulent mixing zone. However, it is important to appreciate that, as we have only considered one particular combination of control parameters ($\mathrm{Pr}$, $\mathrm{Re}_{0}$, $\mathrm{Ri}_{0}$, and $t_{\textrm{r}}$), any observed relationship between the various length scales (particularly with reference to the final depth of the turbulent mixing zone) should be treated with caution.
As shown in section 3.2, for our particular choice of parameters at least, the depth of the turbulent mixing zone increases as the domain becomes horizontally more extended. At one level, this can be explained by an increased number of degrees of freedom of the underlying dynamical system, leading to a higher potential complexity, and thus more vigorous turbulence, causing both a stronger macroscopic stirring and a stronger microscopic mixing. However, there also appears to be a ‘tuning’ of the overall strength of the stratification towards an equilibrium state. Mixing actually increases the ‘strength’ of the stratification (as quantified by either the bulk Richardson number $\mathrm{Ri}$ or appropriate averages of the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$) through deepening the shear layer depth $\langle d\rangle_{t}$ and the buoyancy interface thickness $\langle\delta\rangle_{t}$, see again figure 4. For the turbulence to remain sustained, this deepening cannot continue without limit since turbulence cannot be sustained when the stratification becomes too strong. The particular observation that the (mixing zone) average of $\mathrm{Ri}_{\textrm{g}}$ is approximately $0.15$ is highly reminiscent of the results reported by Portwood et al. (2019). They used a control strategy (effectively through modulating gravity) to identify equilibrium turbulent states in linearly stratified flows driven by constant vertical shear. Although the flow considered here is different, the key dynamics appears to be similar: the forced flow adjusts until the stratification is as strong as possible to still allow for vigorous turbulence which is able to stir and mix the buoyancy field essentially as a passive scalar. This is underlined by the fact that $\langle d\rangle_{t}\approx\langle\delta\rangle_{t}$, and so the mixing of buoyancy is effectively slaved to the mixing of momentum by the turbulence.

### v2

@@ 3.4 Anisotropy of emergent large-scale dynamics
While the previous section 3.3 has made clear that vertical aspects of the flow dynamics converge for horizontally highly extended domains (only), section 3.1 has demonstrated that the considered forced stratified shear flows also develop certain characteristic horizontal structures such as the primary Kelvin-Helmholtz billows. Therefore, it is appropriate to investigate whether there are at later times, when such early billows have broken down, emergent horizontally-aligned structures in the statistically steady flow as well.
@@
Figure 6 shows both the early and final emergent dynamics present in our most extended square domain of $L_{\mathrm{h}}=512$. Panels (a – d, i – l) depict the entire horizontal cross-section at midplane, $\Phi\left(z=0\right)$, whereas panels (e – h, m – p) depict an associated vertical slice at $\Phi\left(y=0\right)$. Note that the vertical slices of $b$ from panels (h, p) are reminiscent of figure 2 (c, e) despite the domain now being $16$ times as large. A video of the evolution of this flow – from $t=0$ to $t=t_{\textrm{evo}}$ – is provided as supplemental material.
At the earlier time, as shown in panel (h), the flow is – analogously to the simulation discussed in section 3.1 and shown in figure 2 – clearly associated with the growth of the initial KHI and the subsequent non-linear formation of overturning billows. Interestingly, as becomes clear via the associated horizontal slice in panel (d), although the initial overturning billows may extend across the entire extended spanwise direction, knots, tubes and defects between these rolls may introduce imperfections to this otherwise regular pattern. This effect has also previously been observed experimentally (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and numerically (Fritts et al., 2022a, b) for (in the spanwise direction) sufficiently wide domains. However, these early aspects of the flow dynamics appear (at least superficially) to be absent or smeared out by the sustained turbulence at late times as shown in panels (l, p).
@@
Independently of the specific point in time, this comparison of vertical slices with the corresponding horizontal midplanes demonstrates the co-existence of apparently quasi-regular dynamics in both the streamwise and spanwise directions. However, these visualisations suggest that the emergent large-scale dynamics exhibits a strong scale separation and, thus, anisotropic associated length scales $\Lambda$. More specifically, at least qualitatively the horizontal structures seem to be much, much larger than the shear layer half-depth, i.e. $\left\{d,\delta\right\}\ll\Lambda_{\textrm{h}}$.
@@
of various quantities. Here, $\hat{\Phi}$ represent the Fourier coefficients corresponding to a transform in either the streamwise or spanwise direction, the asterisk denotes the complex conjugate, and $k_{x}$ and $k_{y}$ are the associated streamwise and spanwise components of the wave vector, respectively. In order to allow for a direct comparability of these spectra with their corresponding terms in the kinetic energy or buoyancy variance equation via Parseval’s theorem, the coefficient $C$ depends on $\Phi_{1}$ and $\Phi_{2}$ as follows: $C=1/2$ for $\Phi_{1}=\Phi_{2}\in\left\{u_{x},u_{y},u_{z},b\right\}$, $C=\mathrm{Ri}_{0}$ for $\Phi_{1}=u_{z}\wedge\Phi_{2}=b$ (or vice versa), $C=2/\mathrm{Re}_{0}$ for $\Phi_{1}=\Phi_{2}=\bm{S}$, or $C=1/\left(\mathrm{Re}_{0}\mathrm{Pr}\right)$ for $\Phi_{1}=\Phi_{2}=\nabla b$. We summarise the key results in figure 7.
@@
Figure 7: Magnitude quantification of emergent large-scale dynamics. (a) Emergent flow structures offer pronounced spectral peaks $\hat{\lambda}$ in the streamwise directions for all variables $\left\{u_{x},u_{y},u_{z},b\right\}$. In contrast, (b) only the streamwise velocity $u_{x}$ exhibits a pronounced peak in spanwise direction. (c) A systematic comparison of characteristic extensions of emergent flow structures along the streamwise, spanwise, and vertical directions highlights a strong anisotropy of the large-scale dynamics. Note that while the coloured lines in panels (a, b) correspond to our largest square domain with $L_{\mathrm{h}}=512$, grey lines are extracted from even more extended but non-square domains ($L_{x}=2048$ and $L_{y}=512$ or $L_{y}=2048$ and $L_{x}=512$). Moreover, as the dash-dotted line $L_{x}=L_{y}$ in panel (c) demonstrates that flow structures may clearly be limited or affected by horizontally insufficiently extended domains, only horizontal extents $L_{\mathrm{h}}\gtrsim 256$ are large enough to resolve the most extended flow structures in the horizontal direction.
@@
In figure 7 (a), we plot the streamwise energy spectra of $\left\{u_{x},u_{y},u_{z},b\right\}$, subject to appropriate spatio-temporal averages. Remarkably, we find that all these flow fields exhibit pronounced spectral peaks at streamwise wave numbers $\hat{k}_{x}\sim\mathcal{O}\left(10^{-2}\right)$. This implies that a certain (narrow range of) streamwise wavelength(s) $\hat{\lambda}_{x}=2\pi/\hat{k}_{x}$ is particularly energetic. In other words, these pronounced spectral peaks establish the existence of flow structures with a characteristic streamwise length scale $\Lambda_{x}\sim\mathcal{O}\left(10^{2}\right)$ in all of these flow fields, thus implying that there is a preferred length scale for the emergent streamwise self-organisation of the large-scale dynamics. We propose a potential dynamical origin of this characteristic scale $\Lambda_{x}$ in section 4.
@@
As shown in figure 7 (b), this emergent property of the streamwise spectra is in clear contrast to the behaviour of the spanwise spectra. On the one hand, we find a similarly pronounced spectral peak only for the $x$-component of the velocity field in the spanwise direction. Together with $\hat{k}_{y}\sim\mathcal{O}\left(10^{-1}\right)$, this implies again the existence of a characteristic spanwise length scale $\Lambda_{y}\sim\mathcal{O}\left(10^{1}\right)$. On the other hand, the other flow fields do not exhibit such a peak but rather flatten out at small $k_{y}$, demonstrating that there is no preferred characteristic length scale for the spanwise self-organisation of the large-scale dynamics in $\left\{u_{y},u_{z},b\right\}$. We remark that even though $E_{u_{y}u_{y}}(k_{y,\textrm{min}})<E_{u_{y}u_{y}}(\hat{k}_{y})$, this difference is smaller than a factor of two and so we do not consider the associated maximum to be a ‘pronounced’ spectral peak.
@@
It is natural to ask whether there could be additional spectral peaks at even smaller wave numbers, i.e., whether the large-scale dynamics might still be biased or limited by the finite domain of $L_{\mathrm{h}}=512$. For this reason, we have conducted two additional simulations which increase either the streamwise or the spanwise extent of the domain by another factor of four. This results in $L_{x}\times L_{y}=2048\times 512$ or $512\times 2048$, respectively. We include the associated energy spectra in figure 7 (a, b) as grey solid lines. We find no evidence of such additional spectral peaks. Furthermore, as both the location and amplitude of the peaks from these spectra coincide with the ones from our largest square domain, we are confident that there is strong evidence that these spectral peaks are characteristic of this particular flow and, crucially, independent of the horizontal extent of the domain. This implies that the large-scale dynamics is (at such large $L_{x}$ and $L_{y}$) indeed governed by mechanisms intrinsic to the flow. This is supported by appendix B, where we show that the energy spectra derived from smaller (yet sufficiently large) domains are also shown to converge with the present ones from the largest domains.
@@
$L_{x}\times L_{y}\thickspace\thickspace$ $E_{u_{x}u_{x}}$ $E_{u_{y}u_{y}}$ $E_{u_{z}u_{z}}$ $E_{bb}$ $E_{u_{z}b}$ $E_{u_{x}u_{x}}$ $2d$ $2\delta$
$16^{2}$ $0$ $\pm\hphantom{0}0 16±016\pm\hphantom{0}$0 $16\pm\hphantom{0}0$ $\pm\hphantom{0}0 16±016\pm\hphantom{0}$0 $15\pm\hphantom{0}3$ $5.3\pm 0.3 0$$5.6\pm 0.3$
$32^{2}$ $0$ $\pm\hphantom{0}0 32±032\pm\hphantom{0}$0 $32\pm\hphantom{0}0$ $\pm\hphantom{0}0 32±032\pm\hphantom{0}$0 $19\pm\hphantom{0}6$ $0.7\pm 0.7 11$$1.1\pm 0.7$
$64^{2}$ $0$ $\pm\hphantom{0}0 64±064\pm\hphantom{0}$0 $64\pm\hphantom{0}2$ $\pm\hphantom{0}0 64±064\pm\hphantom{0}$0 $33\pm 14$ $15.1\pm 0.9$ $15.5\pm 0.7$
@@
$128^{2}$ $0$ $128\pm\hphantom{0}0$ $89\pm 32$ $77\pm 27$ $108\pm 30$ $89\pm 32$ $52\pm 24$ $16.1\pm 0.8$ $16.8\pm 0.9$
$256^{2}$ $0$ $117\pm 18$ $84\pm 13$ $74\pm 12$ $092\pm 2$0 $80\pm 12$ $52\pm 27$ $15.8\pm 0.4$ $16.4\pm 0.3$
[…]
@@
Table 2: Emergent dynamical length scales. This table quantifies the streamwise, spanwise, and vertical extent of the emergent large-scale dynamics via the wavelength associated with the spectral peaks and total depths of the shear layer or density interface. We list values of the temporal mean and standard deviation, the latter of which might be significant due to the discrete nature of wave numbers. Unreliable values, i.e. those clearly affected by the (insufficiently extended) finite domain, are displayed in grey.
@@
Quantifying characteristic length scales associated with the emergent large-scale dynamics, we extract the wavelengths corresponding to these pronounced spectral peaks (or these spectra’s global maxima) for each simulation and summarise them in table 2. Additionally, figure 7 (c) shows this dependence on $L_{\mathrm{h}}$. For small $L_{\mathrm{h}}$, as highlighted by the dash-dotted line, these extracted horizontal length scales $\Lambda_{\left\{x,y\right\}}$ are clearly biased by the extent of the domain. However, this changes once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$ with the critical horizontal extent $L_{\mathrm{h,crit}}$ depending on the solution field (and, thus, the associated final characteristic length scale). While most of the peaks are already properly represented given $L_{\mathrm{h,crit}}=96$, some of them – such as $\hat{\lambda}_{x}\left(E_{u_{x}u_{x}}\right)$ – may require $L_{\mathrm{h}}=256$ to be fully resolved. Hence, for the set of control parameters considered here, a full convergence of the characteristic length scales associated with the large-scale dynamics is reached for $L_{\mathrm{h}}\gtrsim 256$ only. This convergence of finite horizontal scales is, at least qualitatively, consistent with the convergence of the finite vertical depth of the shear layer as addressed in the previous sub-sections.
@@
In addition to the characteristic streamwise and spanwise length scales associated with the large-scale dynamics, we include their characteristic vertical extent via the total shear layer depth in figure 7 (c). Note that since $d\approx\delta$ or $R\approx 1$, these data points equivalently show the total buoyancy interface depth. Comparing the characteristic length scales in the three spatial directions (as denoted by different symbol types), there is a clearly apparent anisotropy of the large-scale dynamics. While vertical scales $\Lambda_{z}\approx 16$ (circles) and spanwise scales $\Lambda_{y}\approx 50$ (upright triangles), streamwise scales (sideways triangles) may extend up to $\Lambda_{x}\approx 115$. Consequently, we find a hierarchy $\Lambda_{z}<\Lambda_{y}<\Lambda_{\textrm{x}}$ that spans across one order of magnitude.
@@
Having extracted these characteristic scales of large-scale dynamics, there are two particularly striking observations. Firstly, while $u_{z}$ and $b$ are related to the vertical transport of buoyancy via equation (15), their individual scales $\hat{\lambda}_{x}$ in $E_{u_{z}u_{z}}$ and $E_{bb}$ differ significantly. Secondly, the total depth of the mixing zone, representing the characteristic vertical extent $\Lambda_{z}$, actually converges at smaller domain extents $L_{\mathrm{h}}$ compared with $\Lambda_{x}$ from $E_{bb}$. These two observations suggest the natural question as to why or how can the vertical characteristics (in both the velocity and buoyancy field) converge before the domain is sufficiently large to allow the separate flow fields to converge horizontally?
@@
Figure 8: Buoyancy exchange and dissipation. (a) The buoyancy flux (i.e. the vertical buoyancy advection) across the midplane – $B\left(x,y,z=0,t=t_{\textrm{evo}}\right)$ for $L_{\mathrm{h}}=512$ – generally stabilises the configuration. (b) The associated streamwise co-spectrum establishes the existence of a characteristic corresponding scale via a pronounced peak, which is in contrast to the spanwise direction. (c) Most of the dissipation, however, is associated with smaller scales similar to the mixing zone depth. In panels (b, c), coloured and grey spectra correspond – similar to figure 7 (a, b) – to the largest square and non-square domains, respectively.
@@
4 Discussion